Thank you so much for your time.
log2(log2(x))=1


Answer by stanbon(48554)

log2(log2(x))=1
Working from the outside in get:
Write the outer log in exponential form, as follows:
2^1=log2(x)
Now write this in exponential form, as follows:
Then 2^2=x
x=4
Cheers,
Stan H.